1. A train, starting at noon, travels north at 40 mph. Another train, starting from the same point at 2pm, travels east at 30 mph. Find to the nearest mile/hr how fast the 2 trains are separating at 3pm.

Answer by mananth(16949)

Train A travels 120 miles in 3 hours
Train B travels 30 miles in 1 hour
..
they are at right angles. The distance between them is the hypotenuse
..
120^2+30^2= D^2
14400+900=D^2
15300 =D^2
D= 123.7 miles